# MOHPER: Multi-objective Hyperparameter Optimization Framework for E-commerce Retrieval System

### v1

@@ Abstract.
E-commerce search optimization has evolved to include a wider range of metrics that reflect user engagement and business objectives. Modern search frameworks now incorporate advanced quality features, such as sales counts and document-query relevance, to better align search results with these goals. Traditional methods typically focus on click-through rate (CTR) as a measure of engagement or relevance, but this can miss true purchase intent, creating a gap between user interest and actual conversions. Joint training with the click-through conversion rate (CTCVR) has become essential for understanding buying behavior, although its sparsity poses challenges for reliable optimization. This study presents MOHPER, a Multi-Objective Hyperparameter Optimization framework for E-commerce Retrieval systems. Utilizing Bayesian optimization and sampling, it jointly optimizes both CTR, CTCVR, and relevant objectives, focusing on engagement and conversion of the users. In addition, to improve the selection of the best configuration from multi-objective optimization, we suggest advanced methods for hyperparameter selection, including a meta-configuration voting strategy and a cumulative training approach that leverages prior optimal configurations, to improve speeds of training and efficiency. Currently deployed in a live setting, our proposed framework substantiates its practical efficacy in achieving a balanced optimization that aligns with both user satisfaction and revenue goals.
@@ 1. Introduction
Initially, search systems relied heavily on keyword-matching algorithms such as TF-IDF, and BM25, which were limited to capturing direct lexical overlap between search keywords and product listings (Ramos et al., 2003; Robertson et al., 2009). However, the recent breakthrough of large language models (LLMs) or neural multi-modal retrieval models makes systems process dense retrievals, by measuring semantic similarities of query-document vectors, as introduced by several studies (Arevalillo-Herráez et al., 2013; Li et al., 2023a; Bhagdev et al., 2008). Although most previous studies focus on relevance scores of the hybrid retrieval using both BM25 and dense retrieval, these similarity scores do not imply the preferences, interests, or engagements of users, which are crucial for e-commerce, generally available by interpreting user interaction logs. To leverage user activity log data, such as selling, clicks, etc. from actual e-commerce service, e-commerce search optimization has evolved markedly, moving from traditional keyword-based retrieval methods to advanced frameworks incorporating relevance, engagement, and an array of business-critical metrics.
@@
Historically, search optimization has centered on maximizing click-through rate (CTR) as an engagement metric, but CTR alone is often insufficient to predict actual conversions, a critical focus in e-commerce. To address this, click-through conversion rate (CTCVR) has become essential for capturing purchase intent. However, training solely on CTCVR presents challenges due to sample selection bias and data sparsity, where training data derived from clicks does not generalize well to the broader space of all impressions, and conversion data remains limited, risking overfitting to a small subset of items (Zhang et al., 2022; Ma et al., 2018). Traditional retrieval or ranking methods, including BM25 (Robertson et al., 2009), dense vector retrievals (Wang et al., 2021; Li et al., 2023a, b), and learning-to-rank (LTR) (Burges et al., 2005), typically optimize for single objectives, which can result in misalignment between relevance, user engagement, and conversion. This single-objective focus limits adaptation to rapidly changing e-commerce trends and degrades performance in online environments (Karmaker Santu et al., 2017).
In response, multi-objective strategies are gaining traction in e-commerce, aiming to optimize both CTR and CTCVR to align engagement with purchase intent. This study presents MOHPER, a multi-objective hyperparameter optimization (MOHPO) framework designed for e-commerce search. Using Bayesian optimization as (Li and Kanoulas, 2018) and advanced sampling techniques, MOHPER jointly optimizes CTR and CTCVR, offering a balanced solution to enhance both engagement and conversions. We systematically evaluate sampling strategies within MOHPER to ensure robust multi-metric optimization, addressing the demand for adaptive, multi-objective solutions in e-commerce. The main contributions of this paper include:
@@
    introduction of a novel hyperparameter optimization (HPO) framework specifically tailored for e-commerce search services, focusing on objectives such as CTR and CTCVR essential indicators of user engagement and purchase intent. This is the first application of a multi-objective HPO framework in e-commerce retrieval and ranking tasks.
@@
    suggestion for advanced methodologies for hyperparameter selection, including a meta-configuration voting strategy and a cumulative training approach that leverages prior optimal configurations.
@@
The notable performance improvements indicate that models optimized via MOHPER have been successfully adopted as foundational models on the OHouse platform, showcasing the practical effectiveness of our approach in real-world e-commerce applications.
2. Proposed Framework: MOHPER
In this section, we present the proposed MOHPER framework. We begin by defining the multi-objective problem within retrieval systems and provide an overview of the HPO algorithm. Subsequently, we identify the relevant objectives for joint optimization in the e-commerce setting. To enhance performance during the meta-evaluation stage and to select appropriate configurations in the multi-objective optimization context, we introduce two methods: meta-configuration voting and cumulative hyperparameter stacking. The proposed methods are illustrated in Figure 1.
2.1. Multi-objective HPO in Retrieval System
For MOHPO, the objective is to identify a set of HPs that optimizes the performance across multiple objectives. These objectives typically conflict, meaning improving one objective may negatively impact another. The goal of HPO is to find an optimal configuration of HPs that balances these objectives. Similarly, in the context of a retrieval system, MOHPO aims to improve the system’s ability to return relevant results for a given set of queries. The objectives generally focus on relevance, ranking quality, and overall system performance. Some common performance metrics for retrieval systems can be either normalized discounted cumulative gain (nDCG), mean average precision (mAP), precision, or recall (Järvelin and Kekäläinen, 2002; Zhu, 2004) while optimizing different labels indicating relevance level, user engagement, or freshness of items, etc.
Figure 2. Illustration of transform function to modulate query request for a search engine as the change of config hyperparameter.
@@ 2.1. Multi-objective HPO in Retrieval System
Formally, the goal is to identify a set of HPs ($h\in\mathcal{H}$) that maximizes the performance of a retrieval engine across multiple objectives $O=\{O_{1},O_{2},\dots,O_{M}\}$ simultaneously. Given an observation dataset $\mathcal{D}_{0}=\emptyset$ and a set of ground truth labels $Y=\{y_{1,(\cdot)},y_{2,(\cdot)},\dots,y_{m,(\cdot)}\}$ for each objective, the optimization seeks to sample and evaluate HP configurations to approximate optimal performance iteratively. In other words, the search system is evaluated by selecting an HP configuration $h_{t}$ from the search space $\mathcal{H}$ at each trial $t=1,2,\dots,N$. During optimization progress, various sampler methods, such as random search, grid search, or Bayesian optimization (e.g. tree-structured Parzen estimator (TPE) or Gaussian process (GP), etc.), can be employed to sample HP configurations. At each trial, $t$, an HP configuration $h_{t}$ is sampled, and the retrieval system is evaluated on a set of queries. The configuration $h_{t}$ influences the retrieval process through transform functions that adapt the retrieval engine’s behavior based on the sampled HPs. This could involve altering ranking strategies, modifying relevance weights, or adjusting search heuristics (see Figure 2). Specifically, the transform function can be expressed as $T(h_{t},q)$, which represents the transformed retrieval process based on the HP configuration $h_{t}$ applied to query $q$. Depending on the sampled HPs, this transformation could modify how items are ranked or filtered for search engines. For each query $q$, the system produces predictions (ranking of items), and the metrics of relevance/quality for these items are evaluated using objectives, respectively. The metric for objective $m$ can be computed based on the predicted rankings $\hat{y}_{h_{t},q,(\cdot)}$ and compared to the ground truth labels $y_{m,q,(\cdot)}$, formulated as $z_{t,m,q}=O_{k}(y_{m,q,(\cdot)},\hat{y}_{h_{t},q,(\cdot)})$. The performance evaluations across all queries are then aggregated to produce a performance metric for each objective. The aggregated metrics are used to update the optimization process, guiding the search for better hyperparameter configurations. This process continues iteratively until convergence, where the goal is to select a hyperparameter configuration that optimizes the trade-off between the multiple objectives. The overall summarized algorithm with TPE sampler is summarized in Algorithm 1.
@@
$l(\mathbf{h})=p(\mathbf{h}\mid z_{t,m}\leq\gamma_{t,m}),\quad g(\mathbf{h})=p(\mathbf{h}\mid z_{t,m}>\gamma_{t,m})$
@@
$\text{EI}(\mathbf{h})=\prod_{m=1}^{M}\frac{l(\mathbf{h})}{g(\mathbf{h})},\quad h_{t}=\arg\max_{\mathbf{h}\in\mathcal{H}}\text{EI}(\mathbf{h})$
@@ 2.2. Joint Optimization in E-Commerce Context
    Click-Through Conversion Rate (CTCVR): =#$=\frac{\#purchase}{\#imp}$
@@
    Click-Through Rate (CTR) =#$=\frac{\#click}{\#imp}$
@@
    Click-Through Cart-Adding Rate (CTCAR): =#$=\frac{\#cart}{\#imp}$
Here, #$\#imp$ represents the number of item impressions in retrievals. CTCVR is the primary objective, with CTR, and CTCAR serving as auxiliary objectives to mitigate selection bias and sparsity (Figure 3) (Ma et al., 2018; Karmaker Santu et al., 2017). Specifically, For any query $q$ and trial $t$, performance metrics $z_{t,q,(\cdot)}$ (e.g. precision@K, nDCG@K, mAP@K, and recall@K) are computed when ground truth labels $y_{q,(\cdot)}$ are available and the top-K items are retrieved ($\hat{y}_{h_{t},q,(\cdot)}$). Labels for these items are derived from objective-specific statistics (e.g. #$\#imp_{q,i}$, #$\#click_{q,i}$, #$\#purchase_{q,i}$), with thresholding for positive/negative scoring at the corresponding objective. Also, while using naive statistical rates is available, filtering based on impression counts and Bayesian smoothing for rare events can further refine these scores (Wang et al., 2011). Following metric calculation $z_{t,m,q}$, metrics are aggregated across queries, and MOHPO proceeds with multi-objective tuning to enhance engagement, conversion efficiency, and purchase intent, advancing retrieval performance in e-commerce.
@@ 2.3. Hyperparameter Extraction Stage
Upon completing HPO training, the framework selects a diverse set of candidate hyperparameter configurations based on their performance across multiple objectives, such as training targets or a weighted metric sum (see the middle stage of Figure 1). These top configurations are then stored for further meta-level evaluation to estimate online performance and for use in subsequent cumulative training (see subsection 2.5).
@@ 2.4. Meta Stage Evaluation and HP selection
Selecting hyperparameter configurations based solely on training metrics can lead to overfitting, as HPO is often biased toward training data and may not generalize well to production due to sample selection bias and data sparsity issues. To address this, the meta-level evaluation assesses top-performing HP configurations on unseen queries and future periods using time-based and query-wise data splits, ensuring no overlap with training data. Each HP is evaluated for robustness across multiple metrics, and we employ a meta-configuration voting strategy (Chan and Stolfo, 1995) to prioritize HPs that demonstrate consistent, balanced performance across objectives. This approach synthesizes the strengths of various configurations, ensuring the selection of parameters that generalize effectively.
@@ 2.5. Cumulative Training
While the baseline algorithm (Sections 3.1–3.4) can operate independently, incorporating prior optimal hyperparameters into each successive MOHPO iteration enhances both performance and stability. Our cumulative training approach initializes each new training cycle with a dataset $\mathcal{D}_{0}=\{h_{prev\ A},z_{prev\ A,1},z_{prev\ A,2},\cdots\}$ in cumulative training, instead of initializing observation dataset $\mathcal{D}_{0}=\emptyset$. This setup allows the sampler to concentrate on unexplored regions, minimizing redundant samples and promoting exploration within the configuration space.
This cumulative method effectively prioritizes configurations with consistently strong performance across objectives, reducing computational redundancy and adapting dynamically to changes in user behavior and market trends. By establishing an adaptive feedback mechanism, cumulative training fortifies the framework’s robustness in evolving e-commerce environments.
@@ 3. Experiments
We conduct extensive experiments to evaluate the effectiveness and the performance of the MOHPO framework and its components and the following research questions (RQs) are answered:
@@ 3.1. Datasets and Experimental Setups
Table 1. The basic statistics of datasets using train/test.
Dataset # Queries # Impressions # Clicks # Conversions
SERP Sampled Train 1,000 367,544,388 12,021,645 1,214,896
SERP Sampled Test 5,000 377,367,483 11,978,753 1,230,871
[…]
CPLP Test 1,255 106,484,599 2,588,947 228,671
  • SERP refers to the search engine result page, employing keywords as queries, and CPLP means the product list page, utilizing category IDs for filtering products. The dataset comprises a subset of production data for baseline comparisons. For online training, the other queries, including tail queries, are integrated into the offline dataset to ensure comprehensive training/evaluation (5,000 for training and 30,000 for testing).
@@ 3.1.2. Retrieval Engine and Indexing Setting
For our experiments, we employ Elasticsearch (ES) [^2] https://github.com/elastic/elasticsearch, specifically version 8.14 for offline training and evaluation tasks and version 8.8 for online experiments. Built upon the Apache Lucene library (McCandless et al., 2010; Białecki et al., 2012), ES is a real-time, scalable, distributed, and multi-tenant search engine supporting both textual and vector search capabilities, widely adopted by several e-commerce websites. Textual features are indexed via inverted indices (Salton et al., 1983), while vector features leverage hierarchical navigable small-world (HNSW) indexing (Malkov and Yashunin, 2018). For experimental consistency, we maintain separate ES instances for offline evaluation and production use, with partial item indices from production duplicated for experimental use. These indices are configured on 3 primary shards with 1 replica each, yielding a total of 6 shards.
Given the multiple retrieval operations required per training trial in HPO, we use the multi-search API [^3] https://www.elastic.co/guide/en/elasticsearch/reference/current/search-multi-search.html to execute concurrent searches within a single request, similar to a mini-batch bulk request approach. This concurrent processing is essential for the HPO process, where search results are repeatedly requested to generate predictions based on various hyperparameter configurations, significantly improving training efficiency through parallelization.

### v2

@@ Abstract.
E-commerce search optimization has evolved to include a wider range of metrics that reflect user engagement and business objectives. Modern search frameworks now incorporate advanced quality features, such as sales counts and document-query relevance, to better align search results with these goals. Traditional methods typically focus on click-through rate (CTR) as a measure of engagement or relevance, but this can miss true purchase intent, creating a gap between user interest and actual conversions. Joint training with the click-through conversion rate (CTCVR) has become essential for understanding buying behavior, although its sparsity poses challenges for reliable optimization. This study presents MOHPER, a Multi-Objective Hyperparameter Optimization framework for E-commerce Retrieval systems. Utilizing Bayesian optimization and sampling, it jointly optimizes both CTR, CTCVR, and relevant objectives, focusing on engagement and conversion of the users. In addition, to improve the selection of the best configuration from multi-objective optimization, we suggest advanced methods for hyperparameter selection, including a meta-configuration voting strategy and a cumulative training approach that leverages prior optimal configurations, to improve speed of training and efficiency. Currently deployed in a live setting, our proposed framework substantiates its practical efficacy in achieving a balanced optimization that aligns with both user satisfaction and revenue goals.
@@ 1. Introduction
Traditional search systems often rely on lexical retrieval techniques such as TF-IDF and BM25 (Ramos et al., 2003; Robertson et al., 2009), which match keywords between user queries and product documents. More recently, neural retrieval models based on large language models (LLMs) (Arevalillo-Herráez et al., 2013; Li et al., 2023a; Bhagdev et al., 2008) and multimodal embeddings have improved semantic alignment (Radford et al., 2021), enabling richer query-document representations. Although most previous studies focus on relevance scores of the hybrid retrieval using both BM25 and dense retrieval, these similarity scores do not imply the preferences, interests, or engagements of users, which are crucial for e-commerce, generally available by interpreting user interaction logs. Click-based optimization, especially CTR, has become a standard proxy for user preference. Yet, CTR is an imperfect signal: it overestimates interest and underrepresents conversions. This mismatch motivates the use of more direct metrics such as click-through-conversion rate (CTCVR), which captures actual purchase intent. However, conversion is inherently sparse and biased, since it only exists for a subset of clicked items, risking overfitting to a small subset of items (Zhang et al., 2022; Ma et al., 2018). Traditional retrieval or ranking methods, including BM25 (Robertson et al., 2009), dense vector retrievals (Wang et al., 2021; Li et al., 2023a, b), and learning-to-rank (LTR) (Burges et al., 2005) typically optimizes for single objectives, which can result in misalignment between relevance, user engagement, and conversion. This single-objective focus limits adaptation to rapidly changing e-commerce trends and degrades performance in online environments (Karmaker Santu et al., 2017).
@@
In response, multi-objective strategies are gaining traction in e-commerce, aiming to optimize both CTR and CTCVR to align engagement with purchase intent. This study presents MOHPER, a multi-objective hyperparameter optimization (MOHPO) framework designed for e-commerce search. Using Bayesian optimization as (Li and Kanoulas, 2018) and advanced sampling techniques, MOHPER jointly optimizes CTR and CTCVR, offering a balanced solution to enhance both engagement and conversions in selecting optimal hyperparameters (HPs). We systematically evaluate sampling strategies within MOHPER to ensure robust multi-metric optimization, addressing the demand for adaptive, multi-objective solutions in e-commerce. The main contributions of this paper include:
@@
    introduction of a novel hyperparameter optimization (HPO) framework specifically tailored for e-commerce search services, focusing on objectives such as CTR and CTCVR, essential indicators of user engagement and purchase intent. This is the first application of a multi-objective HPO framework in e-commerce retrieval and ranking tasks.
@@
    suggestion for advanced methodologies for HP selection, including a meta-configuration voting strategy and a cumulative training that leverages prior optimal configurations.
@@
2. MOHPER: Problem Formulation and Design
In this section, we present the proposed MOHPER framework. We begin by formulating the multi-objective problem within retrieval systems and provide an overview of the HPO algorithm. Subsequently, we identify the relevant objectives for joint optimization in the e-commerce setting. To enhance performance during the meta-evaluation stage and to select appropriate configurations in the multi-objective optimization context, we introduce two methods: meta-configuration voting and cumulative HP stacking. The proposed methods are illustrated in Figure 1.
2.1. Multi-Objective HPO in Retrieval Systems
Formally, the goal is to identify a set of HPs ($h\in\mathcal{H}$) that maximizes the performance of a retrieval engine across multiple objectives $O=\{O_{1},O_{2},\dots,O_{M}\}$ simultaneously. Given an observation dataset $\mathcal{D}_{0}=\emptyset$ and a set of ground truth labels $Y=\{y_{1,(\cdot)},y_{2,(\cdot)},\dots,y_{m,(\cdot)}\}$ for each objective, the optimization seeks to sample and evaluate HP configurations to approximate optimal performance iteratively. In other words, the search system is evaluated by selecting an HP configuration $h_{t}$ from the search space $\mathcal{H}$ at each trial $t=1,2,\dots,N$. During optimization progress, various sampler methods, such as random search, grid search, or Bayesian optimization (e.g., tree-structured Parzen estimator (TPE) or Gaussian process (GP), etc.), can be employed to sample HP configurations. At each trial, $t$, an HP configuration $h_{t}$ is sampled, and the retrieval system is evaluated on a set of queries. The configuration $h_{t}$ influences the retrieval process through transform functions that adapt the retrieval engine’s behavior based on the sampled HPs. This could involve altering ranking strategies, modifying relevance weights, or adjusting search heuristics (see Figure 2). Specifically, the transform function can be expressed as $T(h_{t},q)$, which represents the transformed retrieval process based on the HP configuration $h_{t}$ applied to query $q$. Depending on the sampled HPs, this transformation could modify how items are ranked or filtered for search engines. For each query $q$, the system produces predictions (ranked items), and the relevance/quality metrics for these items are evaluated using objectives, respectively. The metric objective $m$ can be computed based on the predicted rankings $\hat{y}_{h_{t},q,(\cdot)}$ and compared to the ground truth labels $y_{m,q,(\cdot)}$, formulated as $z_{t,m,q}=O_{k}(y_{m,q,(\cdot)},\hat{y}_{h_{t},q,(\cdot)})$. Performance evaluations across all queries are then aggregated to produce a performance metric for each objective. The aggregated metrics are used to update the optimization process, guiding the search for better HP configurations. This process continues iteratively until convergence, where the goal is to select a HP configuration that optimizes the trade-off between the multiple objectives. The overall algorithm with TPE sampler is summarized in Algorithm 1.
Figure 2. Transform function to modulate query request for a search engine as the change of config HP
@@ 2.1. Multi-Objective HPO in Retrieval Systems
$l_{m}(\mathbf{h})=p(\mathbf{h}\mid z_{t,m}\leq\gamma_{t,m}),\quad g_{m}(\mathbf{h})=p(\mathbf{h}\mid z_{t,m}>\gamma_{t,m})$
@@
$\text{EI}(\mathbf{h})=\prod_{m=1}^{M}\frac{l_{m}(\mathbf{h})}{g_{m}(\mathbf{h})},\quad h_{t}=\arg\max_{\mathbf{h}\in\mathcal{H}}\text{EI}(\mathbf{h})$
@@ 2.2. Joint Optimization in E-Commerce Context
    Click-Through Conversion Rate (CTCVR): #$\#\text{purchases}\ /\#\text{imp.}$.
@@
    Click-Through Rate (CTR): #$\#\text{clicks}\ /\ \#\text{imp.}$.
@@
    Click-Through Cart-Adding Rate (CTCAR): #$\#\text{carts}\ /\ \#\text{imp.}$.
[…]
@@ 2.3. Hyperparameter Extraction Stage
Upon completing HPO training, the framework selects a diverse set of candidate HP configurations based on their performance across multiple objectives, such as training targets or a weighted metric sum (see the middle stage of Figure 1). These top configurations are then stored for further meta-level evaluation to estimate online performance and for use in subsequent cumulative training.
@@ 2.4. Meta Stage Evaluation and HP selection
To mitigate overfitting from training data, we introduce a meta-evaluation stage that validates top HP configurations on unseen queries and time-based splits, improving robustness and generalization (Cawley and Talbot, 2010; Shani and Gunawardana, 2011). Each configuration is evaluated across multiple objectives, including training and other service-relevant objectives.
We introduce meta-configuration voting, based on a voting-based selection strategy (Chan and Stolfo, 1995), where each $h\in\mathcal{H}_{\text{top}}$ is scored as:
(1) $\text{Vote}(h)=\sum_{k=1}^{K}\mathbb{1}\{h\in\text{top-}n\text{ on objective }k\},\quad h^{*}=\arg\max_{h}\text{Vote}(h).$
This ensemble-style approach reduces variance (Breiman, 1996) and aligns with stability-based generalization theory (Bousquet and Elisseeff, 2002). Consistent performance across independent validation splits further bounds generalization error via concentration inequalities such as the Chernoff bound (Chernoff, 1952).
@@ 2.5. Cumulative Training
To enhance convergence and sample efficiency, we adopt a cumulative training strategy by reusing top configurations from previous rounds:
(2) $\mathcal{D}_{0}=\{(h,z_{1},\dots,z_{M})\mid z_{m}\geq\gamma_{m}\ \forall m\in\mathcal{O}\}.$
This warm-started optimization avoids redundant exploration and accelerates convergence (Lin, 1992). Cumulative training can be interpreted through an empirical Bayes lens (Morris, 1983), where prior high-performing configurations act as data-driven priors. This improves algorithmic stability and reduces sensitivity to distribution shifts (Bousquet and Elisseeff, 2002). Given the search history $S_{t-1}$, the candidate pool at step $t$ becomes:
(3) $S_{t}=S_{t-1}\cup\{h_{t}\},$
allowing knowledge accumulation over iterations and more efficient exploration of the search space.
@@ 3. Experiments
We conduct extensive experiments to evaluate the effectiveness and the performance of the MOHPO framework and its components, and the following research questions (RQs) are answered:
@@ 3.1.1. Datasets and Split Settings
Table 1. Dataset statistics used in training and evaluation.
Split # Queries # Impressions # Clicks # Conversions
SERP Train 1,000 367.54M 12.02M 1.21M
SERP Test 5,000 377.37M 11.98M 1.23M
CPLP Train 1,379 436.76M 9.58M 0.86M
CPLP Test 1,255 106.48M 2.59M 0.23M
@@ 3.1.2. Retrieval Engine and Indexing Setting
We use Elasticsearch v8.14 for offline experimentation and v8.8 in production. Search indices consist of both sparse and dense embeddings, supporting lexical and semantic retrieval. Dense indexing uses HNSW (Malkov and Yashunin, 2018), and sparse retrieval uses standard inverted indices (Salton et al., 1983).
Each query $q$ is issued to the search engine $R$ via a transform function $T(h,q)$ that modifies the engine’s behavior according to the HP configuration $h$. For efficiency, evaluations are batched using Elasticsearch’s multi-search API [^2] https://www.elastic.co/guide/en/elasticsearch/reference/current/search-multi-search.html to execute concurrent searches within a single request. This concurrent processing is essential for the HPO process, where search results are repeatedly requested to generate predictions based on various HP configurations, significantly improving training efficiency through parallelization.